OpenStudy (anonymous):

if you have four 1's.what is the largest number you can make.(you can use any operator...+,-* etc)

OpenStudy (anonymous):

i think it should just be 4

OpenStudy (anonymous):

it is 4

OpenStudy (anonymous):

sorry..its not...try again

OpenStudy (anonymous):

either 1+1+1+1 or (1+1)(1+1)

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

hint:any mathematical sign you can use

OpenStudy (nowhereman):

Then there is no biggest one, you can always take the factorial: (1+1+1+1)!!!!!!!!!!.. and so on

OpenStudy (anonymous):

no

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

yes,nowhereman,u are corrrect

OpenStudy (anonymous):

answer is--- 11^11

OpenStudy (nowhereman):

So if decimal system is allowed, why is factorial not? That's really arbitrary. If you want to make this a fair question you have to list which operations are allowed.

OpenStudy (anonymous):

i have just used the power.. 11 to the power 11

OpenStudy (anonymous):

use factorial but it will be a smaller no.

OpenStudy (nowhereman):

No it won't: \[(1+1+1+1) = 4, \quad 4! = 24\]\[(1+1+1+1)!! = 24! = 620448401733239439360000 > 285311670611 = 11^{11}\]

OpenStudy (anonymous):

11*11=121

OpenStudy (anonymous):

no fairy.. its 11 to the power 11